OpenStudy (anonymous):

Show your work to prove that the equation v = -4t - 2 correctly represents the function for the data in this table. A t v table shows the following coordinate pairs: negative two comma six, negative one comma two, and zero comma negative two.

jimthompson5910 (jim_thompson5910):

the first row -2 6 means that t = -2 and v = 6

jimthompson5910 (jim_thompson5910):

plug those both in and see if the equation is true

jimthompson5910 (jim_thompson5910):

v = -4t - 2 6 = -4(-2) - 2 6 = 8 - 2 6 = 6 so that proves that t = -2 and v = 6 make the equation true so that row has been confirmed

jimthompson5910 (jim_thompson5910):

do the same to the other rows to fully confirm the table

OpenStudy (anonymous):

should it equal 6 everytime or how do you know if it is correct?

jimthompson5910 (jim_thompson5910):

both sides should be equal when you plug in the t and v values

jimthompson5910 (jim_thompson5910):

it's not going to equal 6 every time

OpenStudy (anonymous):

okay let me do them to see if they are correct

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

I did it and all of them matched thanks so much